OpenStudy (anonymous):

can i please get help with this, i dont know how to solve it Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

there is a formula for sin(2x)

OpenStudy (anonymous):

do you know what it is?

OpenStudy (anonymous):

i dont think so

OpenStudy (anonymous):

sin(2x) = 2sin(x)cos(x)

OpenStudy (anonymous):

Equation: 7 sin2x - 14 sin x + 2 = -5 move 2 to the other side to get?

OpenStudy (mathmale):

@q12157: Very nice start, very helpful. Be certain to let mabitrix find his or her own answer(s), with your guidance.

OpenStudy (anonymous):

7 sin2x -14 sinx = -3

OpenStudy (anonymous):

now plug in the formula for sin(2x) to get?

OpenStudy (anonymous):

also its not -3 its -7

OpenStudy (anonymous):

since you are subtracting 2 on both sides

OpenStudy (anonymous):

7 2sin(x)cos(x) -14 sinx = -7 ?

OpenStudy (anonymous):

7 2sin(x)cos(x) is the same as 14sin(x)cos(x) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

aren't i supposed to make the equation =0?

OpenStudy (anonymous):

question... is it sin(2x) or is it sin^2(x) for the problem...

OpenStudy (anonymous):

if its sin^2(x)...its a totally different process lol

OpenStudy (anonymous):

oh it's sin^2(x)

OpenStudy (anonymous):

lol... 7 sin^2(x) - 14 sin x + 2 = -5 move -5 to other side to get: 7 sin^2(x) - 14 sin x + 7 = 0 Divide everything by 7 to get sin^2(x) - 2sin x + 1 = 0 Correct?

OpenStudy (anonymous):

yea i get it better now

OpenStudy (anonymous):

The trick here is now to take a variable such as B and set that equal to sin(x) so B=sin(x) Thus the equation becomes B^2-2B+1=0

OpenStudy (anonymous):

Do you know how to factor this?

OpenStudy (anonymous):

The equation becomes: (B-1)*(B-1)=0

OpenStudy (anonymous):

Have you learnt multiplicity yet?

OpenStudy (anonymous):

so so i do (sinx - 1)(sinx -1)=0 ?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

so what does sin(x)=?

OpenStudy (anonymous):

but why is it -1 in the parenthesis?

OpenStudy (anonymous):

sinx = 1

OpenStudy (anonymous):

its (sin(x) - 1)(sin(x) -1)=0

OpenStudy (anonymous):

and yep

OpenStudy (anonymous):

since sinx=1 what does x=?

OpenStudy (anonymous):

do i do sin^-1 (1)? and thats equal pi/2?

OpenStudy (anonymous):

yep! except theres a catch...

OpenStudy (anonymous):

The question asks to find "all solutions"

OpenStudy (anonymous):

thus (pi/2)+2pi is also a solution, cause if you go all the way around the circle it still causes sin(x) to equal 1. Thus the real answer is: (pi/2)+2pi*k where k is any positive or negative integer constant

OpenStudy (anonymous):

(yes you'd have to write that full answer down on your hw to receive full credit)

OpenStudy (anonymous):

oh i understand, thank you for your help!